# 12, Proskura st., Kharkov 61085,Ukraine

### v1

@@ Abstract
We consider the linear vector Schrödinger equation subjected to quadratic constraints. We demonstrate that the resulting nonlinear system is closely related to the Ablowitz-Ladik hierarchy and use this fact to derive the $N$-soliton solutions for the discussed model.
@@
In this paper we want to discuss some nonlinear and seemingly integrable model in which the nonlinearity arises from the imposed constraints. We follow the approach developed by Pohlmeyer who considered in [1] the linear wave equation under quadratic constraints. This approach, which has been generalized by various authors, leads to the so-called $\sigma$-models, which play an important role in modern mathematics and physics (see, e.g., [2, 3, 4, 5, 6, 7]).
[…]
@@
where $\boldsymbol{\psi}$ is a two-dimensional complex vector which is a function of two real variables $t$ and $x$, $\boldsymbol{\psi}=\boldsymbol{\psi}(t,x)$, $\boldsymbol{\psi}^{\dagger}$ is its Hermitian conjugate and subscripts denote derivatives with respect to the corresponding variables. The Lagrange multiplier $\lambda(t,x)$ is introduced to met the constraint
@@
$0=i\boldsymbol{\psi}_{t}+\boldsymbol{\psi}_{xx}+\lambda\boldsymbol{\psi},\qquad\lambda=\mathop{\mbox{Im}}\boldsymbol{\psi}^{\dagger}\boldsymbol{\psi}_{t}+\boldsymbol{\psi}^{\dagger}_{x}\boldsymbol{\psi}_{x}.$ (1.3)
The key point of this work is to demonstrate that equations (1.3) can be ‘embedded’ into the Ablowitz-Ladik hierarchy (ALH). In section 2 we show how one can obtain solutions for (1.3) from solutions for the equations of the ALH. Such approach was used in, say, [8, 9, 10] and was shown to be rather useful when one wants to find some particular solutions because the ALH is one of the most well-studied integrable systems. In section 3 we derive the $N$-soliton solutions for our problem by modification of the already known solitons of the ALH.
2 ALH hierarchy.
The ALH was introduced in 1976 in [11] as an infinite number of differential equations, the most famous of which is the discrete nonlinear Schródinger equation.
Later, it has been reformulated as a system of a few functional equations generated by the Miwa shifts applied to the functions of an infinite number of arguments (see [12]). The Miwa shifts, denoted by $\mathbb{E}_{\xi}$, are defined as
@@
$\begin{array}[]{l}0=\tau_{n}\mathbb{E}_{\xi}{\tau_{n}}-\tau_{n-1}\mathbb{E}_{\xi}{\tau_{n+1}}-\rho_{n}\mathbb{E}_{\xi}{\sigma_{n}},\\[5.69054pt] 0=\tau_{n}\mathbb{E}_{\xi}{\sigma_{n}}-\sigma_{n}\mathbb{E}_{\xi}{\tau_{n}}-\xi\tau_{n-1}\mathbb{E}_{\xi}{\sigma_{n+1}},\\[5.69054pt] 0=\rho_{n}\mathbb{E}_{\xi}{\tau_{n}}-\tau_{n}\mathbb{E}_{\xi}{\rho_{n}}-\xi\rho_{n-1}\mathbb{E}_{\xi}{\tau_{n}}.\end{array}\quad n\in\left(-\infty,...,\infty\right).$ (2.4)
@@
Now, we will derive some consequences of (2.4), which we use below to solve our problem. Introducing, for a fixed value of $n$,
@@
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\frac{\sigma_{0}}{\tau_{0}},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}=\frac{\rho_{0}}{\tau_{0}},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}=\frac{\tau_{1}}{\tau_{0}},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=\frac{\tau_{-1}}{\tau_{0}},$ (2.6)
one can show that these functions satisfy
@@
$-i\boldsymbol{r}^{\scriptscriptstyle T}_{t}+\boldsymbol{r}^{\scriptscriptstyle T}_{xx}+\lambda\boldsymbol{r}^{\scriptscriptstyle T}=0$ (2.22)
with $\lambda=2uv$ as well as the identities
$\lambda=-i\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}_{t}-\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}_{xx}=i\boldsymbol{r}^{\scriptscriptstyle T}_{t}\boldsymbol{q}-\boldsymbol{r}^{\scriptscriptstyle T}_{xx}\boldsymbol{q}$ (2.23)
[…]
@@
So, as a by-products, we have obtained solutions for the nonlinear Schrödinger equation (see section 9.2 of [12]).
Till now, the correspondence between the ALH and the model (2.24) was rather general: any solution for (2.4) provide solution for (2.21)–(2.23). However, not all of them can be used to obtain solutions with $\boldsymbol{q}$ and $\boldsymbol{r}$ being relates by, say, $\boldsymbol{r}^{\scriptscriptstyle T}=\boldsymbol{q}^{\dagger}$. One thus need some additional work. Moreover, we have to make some slightly nonstandard steps. The case is that the ‘natural’ involution for the ALH-like equations is $\rho_{n}=\pm\sigma_{n}^{*}$, $\tau_{n}=\tau_{n}^{*}$ where * stands for the complex conjugation. Clearly, such involution does not provide necessary relation between $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}$. Nevertheless, this issue can be resolved. In the next section we construct $N$-soliton solutions for our problem by modifying the already known ones that have been derived earlier for the ALH and discuss the issue of involution in more detail.
@@
The main part of the structure of the soliton solutions are the so-called ‘soliton matrices’, that satisfy the system of Sylvester equations
@@
and that have been repeatedly used in the framework of the Cauchy matrix approach (see, e.g., [13, 14, 15, 16]).
@@
Using the results of the paper [17], one can present functions (2.6) as
@@
It can be shown that, if the Miwa shifts are implemented as
@@
which clearly implies
@@
then the functions (3.5)–(3.8) satisfy (2.7) and (2.8) (see Appendix). Moreover, these functions are solutions of all differential equations of the ALH (equations (2.13)–(2.16) included), provided the variables of the hierarchy, $z_{1},z_{2},...$, are introduced in accordance with the definition (3.10). Returning to our problem, this means that the $(x,t)$-dependency is governed by
@@
where $\langle a_{0}|=\left(a_{01},...,a_{0N}\right)$ and $\langle b_{0}|=\left(b_{01},...,b_{0N}\right)$ are arbitrary constant $N$-rows,
@@
and $\mathsf{D}_{\alpha}$, $\mathsf{D}_{\beta}$, $\mathsf{D}_{a}$, $\mathsf{D}_{b}$ are diagonal matrices with elements $\alpha_{j}$, $\beta_{j}$, $a_{0j}$, $b_{0j}$ ($j=1,...,N$) correspondingly, one can reveal the following structure of the matrices $\mathsf{F}$ and $\mathsf{G}$:
$\mathsf{F}=\mathsf{D}_{\beta}\left(\mathsf{1}+\mathsf{Y}^{*}\mathsf{Y}\right)^{-1}\mathsf{D}_{\beta}^{-1},\qquad\mathsf{G}=\mathsf{D}_{\alpha}\left(\mathsf{1}+\mathsf{Y}\mathsf{Y}^{*}\right)^{-1}\mathsf{D}_{\alpha}^{-1},$ (3.23)
@@
$\mathsf{Y}=\mathsf{C}\mathsf{E},\quad\mathsf{E}=\mathsf{E}_{A}\mathsf{D}_{\beta}\mathsf{D}_{a}.$ (3.24)
@@
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\langle 1|\mathsf{R}^{-1}\mathsf{\Omega}|1\rangle,\quad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}=1+\langle 1|(\mathsf{R}^{*})^{-1}\mathsf{\Omega}\mathsf{Y}^{*}|1\rangle$ (3.25)
@@
$\mathsf{\Omega}=\mathsf{E}\left(\mathsf{1}+\mathsf{Y}^{*}\mathsf{Y}\right)^{-1}$ (3.26)
@@
Looking at (3.25) one can note that $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ are solitons of different type: $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ is a so-called bright soliton, vanishing as $x\to\pm\infty$ (with $t$ being fixed), while $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ is a dark soliton ($\lim_{x\to\pm\infty}|\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}|=1$). However, this does not mean that this is true for any solution, because of the fact that both the equations and constraints are invariant under transformations $\boldsymbol{q}\to\mathsf{U}\boldsymbol{q}$, if $\mathsf{U}^{\dagger}\mathsf{U}=\mathsf{1}$. So, general $N$-soliton solution is a mixture of dark and bright solitons.
@@
$\boldsymbol{\psi}=\mathsf{U}\left(\begin{array}[]{c}\langle\gamma^{*}|\mathsf{\Omega}|1\rangle,\\ 1+\langle\gamma|\mathsf{\Omega}\mathsf{Y}^{*}|1\rangle\end{array}\right)$ (3.28)
where $\mathsf{U}$ is an arbitrary unitary matrix,
$\mathsf{\Omega}=\mathsf{E}\left(\mathsf{1}+\mathsf{Y}^{*}\mathsf{Y}\right)^{-1},\quad\mathsf{Y}=\mathsf{C}\mathsf{E},\quad\mathsf{E}=\exp\left(-i\mathsf{\Theta}\right),$ (3.29)
$\mathsf{C}$ defined in (3.22),
$\mathsf{\Theta}=\mathsf{\Theta}(x,t)=\mathrm{diag}\left(x(L^{*}_{j})^{-1}+t(L^{*}_{j})^{-2}+\delta_{j}\right)_{j=1,..,N}$ (3.30)
and $\langle\gamma|$ is the constant $N$-row,
$\langle\gamma|=\left(1/L_{1},...,1/L_{N}\right),$ (3.31)
@@
solve the Euler–Lagrange equations (1.3). Elements of the matrix $\mathsf{U}$ together with the $3N$ constants $\mathrm{Re}\,L_{j}$, $\mathrm{Im}\,L_{j}$ and $\delta_{j}$ are the arbitrary parameters that determine the properties of the $N$-soliton solution.
@@
A more interesting question is whether we can tackle with the approach of this paper the case on quadratic restrictions other than (1.2), for example ones given by

### v2

# 12, Proskura st., Kharkiv 61085,Ukraine

@@ Abstract
We consider the linear vector Schrödinger equation subjected to quadratic constraints. We demonstrate that the resulting nonlinear system is closely related to the Ablowitz-Ladik hierarchy and use this fact to derive the $N$-soliton solutions for the discussed model. As an example of application of these results we present solitons of some vector nonlinear Schrödinger equation with gradient nonlinearity.
@@
In this paper we want to discuss some nonlinear and seemingly integrable model in which the nonlinearity arises from the imposed constraints. We follow the approach developed by Pohlmeyer who considered in [1] the linear wave equation under quadratic constraints. This approach, which has been generalized by various authors, leads to the so-called $\sigma$-models, which play an important role in modern mathematics and physics (see, e.g., the books [2, 3]).
Here, we would like to find some solutions for the problem when similar constraints are applied to the linear Schrödinger equation. This problem is described by the action $\mathcal{S}=\int dx\,dt\,\mathcal{L}$ with the Lagrangian
@@
where $\boldsymbol{\psi}$ is a two-dimensional complex vector which is a function of two real variables $t$ and $x$, $\boldsymbol{\psi}=\boldsymbol{\psi}(t,x)$, $\boldsymbol{\psi}^{\dagger}$ is its Hermitian conjugate and subscripts denote derivatives with respect to the corresponding variables. The Lagrange multiplier $\lambda(t,x)$ is introduced to meet the constraint
@@
$\begin{array}[]{r}i\boldsymbol{\psi}_{t}+\boldsymbol{\psi}_{xx}+\lambda\boldsymbol{\psi}=0,\\ -i\boldsymbol{\psi}^{\dagger}_{t}+\boldsymbol{\psi}^{\dagger}_{xx}+\lambda\boldsymbol{\psi}^{\dagger}=0\end{array}$ (1.3)
with
$\lambda=\mathop{\mbox{Im}}\boldsymbol{\psi}^{\dagger}\boldsymbol{\psi}_{t}+\boldsymbol{\psi}^{\dagger}_{x}\boldsymbol{\psi}_{x}.$ (1.4)
The key point of this work is to demonstrate that equations (1.3) and (1.4) can be ‘embedded’ into the Ablowitz-Ladik hierarchy (ALH). In section 2 we show how one can obtain solutions for (1.3) from solutions for the equations of the ALH. Such approach was used in, for example, [4, 5] and was shown to be rather useful when one wants to find some particular solutions because the ALH is one of the most well-studied integrable systems. In section 3 we derive the $N$-soliton solutions for our problem by modification of the already known solitons of the ALH. In section 4 we consider an example of possible applications of the obtained results and present solitons of some vector nonlinear Schrödinger equation (NLSE) with gradient nonlinearity.
2 Ablowitz-Ladik hierarchy.
The ALH was introduced in 1976 in [6] as an infinite number of differential equations, the most famous of which is the discrete nonlinear Schrödinger equation.
Later, it has been reformulated as a system of a few functional equations generated by the Miwa shifts applied to the functions of an infinite number of arguments. The Miwa shifts, denoted by $\mathbb{E}_{\xi}$, are defined as
@@
[…]
@@
Now, we will derive some consequences of (2.4), which we use below to solve our problem.
First we introduce, for a fixed value of $n$,
@@
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\frac{\sigma_{0}}{\tau_{0}},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}=\frac{\rho_{0}}{\tau_{0}},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}=\frac{\tau_{1}}{\tau_{0}},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=\frac{\tau_{-1}}{\tau_{0}}.$ (2.6)
The original Ablowitz-Ladik equations were formulated in terms of $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}$ (with the $n$-dependence restored), and the first two equations in (2.6) are the standard way to introduce the tau-functions in order to arrive at bilinear equations (2.4). The second pair of functions, $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}$, usually was not considered (or even introduced) in the framework of the Ablowitz-Ladik equations. However, as we see in what follows, they are a ‘natural’ complement to $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}$ and will play an essential role in this work.
One can show that functions defined in (2.6) satisfy
@@
$i\boldsymbol{r}^{\scriptscriptstyle T}_{t}+\boldsymbol{r}^{\scriptscriptstyle T}_{xx}+\lambda\boldsymbol{r}^{\scriptscriptstyle T}=0$ (2.22)
with
$\lambda=2uv$ (2.23)
as well as the identities
$\lambda=-i\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}_{t}-\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}_{xx}=i\boldsymbol{r}^{\scriptscriptstyle T}_{t}\boldsymbol{q}-\boldsymbol{r}^{\scriptscriptstyle T}_{xx}\boldsymbol{q}.$ (2.24)
It is easy to see that equations (2.21) and (2.22) together with (2.24) are the Euler–Lagrange equations for the action $\mathcal{S}=\int dx\,dt\;\mathcal{L}$ with the Lagrangian
@@
It should be noted that problem with the Lagrangian (2.25) is more general than one with (1.1) and our original problem is a reduction of the former: $\boldsymbol{r}^{\scriptscriptstyle T}=\boldsymbol{q}^{\dagger}$. In physical applications, models involving $\boldsymbol{\psi}$ and $\boldsymbol{\psi}^{\dagger}$ appear more often than similar models with two independent fields, as $\boldsymbol{q}$ and $\boldsymbol{r}$ in our case. However it is a common practice even in the cases of problems with complex fields, like, for example, the NLSE, consider $\boldsymbol{\psi}$ and $\boldsymbol{\psi}^{\dagger}$ as distinct variables because most part (but surely not all) of the calculations depend on the algebraic structure of the equations and not on the involution $\boldsymbol{\psi}\leftrightarrow\boldsymbol{\psi}^{\dagger}$.
@@
(we give a proof of this statement in Appendix A). So, as a by-products, we have obtained solutions for the NLSE.
Till now, the correspondence between the ALH and the model (2.25) was rather general: any solution for (2.4) provides solution for (2.21)–(2.24). However, not all of them can be used to obtain solutions with $\boldsymbol{q}$ and $\boldsymbol{r}$ being related by $\boldsymbol{r}^{\scriptscriptstyle T}=\boldsymbol{q}^{\dagger}$. One thus needs some additional work. Moreover, we have to make some slightly nonstandard steps. The case is that the ‘natural’ involution for the ALH-like equations is $\rho_{n}=\pm\sigma_{n}^{*}$, $\tau_{n}=\tau_{n}^{*}$ where * stands for the complex conjugation. Clearly, such involution does not provide necessary relation between $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}$. Nevertheless, this issue can be resolved. In the next section we construct $N$-soliton solutions for our problem by modifying the already known ones that have been derived earlier for the ALH and discuss the issue of involution in more detail.
@@
The main part of the structure of the soliton solutions are the so-called ‘soliton matrices’, that satisfy the system of the Sylvester equations
@@
and that have been repeatedly used in the framework of the Cauchy matrix approach (see, e.g., chapter 9 of [7]).
@@
The recipe for soliton solutions consists of two parts. The first step is to ‘construct’ functions $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$$\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}$ of the matrices and vectors introduced above. As was mentioned in the Introduction, we use the already known results for the ALH. So, we follow section 2.2 of the paper [8], and present functions (2.6) as
@@
The next step, is the dependence of $\mathsf{A}$, $\mathsf{B}$ etc on the variables of the hierarchy $z_{1}$, $z_{2}$, … (and, in particular, on $x$ and $t$). Again, we do not invent anything new and use the ‘classical’ prescription. The case is that usually, in soliton solutions, the $z_{k}$-dependence appears through various exponential functions like $\exp[-i\phi(\boldsymbol{z})]$ with $\phi(\boldsymbol{z})=\sum_{k=1}^{\infty}c^{k}z_{k}$ (the simplest non-trivial series). The importance of such functions stems from the identity
$\mathbb{E}_{\xi}{\phi}-\phi=i\sum_{k=1}^{\infty}\frac{(c\,\xi)^{k}}{k}=-i\ln(1-c\xi)$ (3.10)
and, as a result, from the fact that $\exp[-i\phi(\boldsymbol{z})]$ is an eigenfunction of the shift operator $\mathbb{E}_{\xi}$:
$\mathbb{E}_{\xi}{e^{-i\phi(\boldsymbol{z})}}=e^{-i\phi(\boldsymbol{z})}/(1-c\xi).$ (3.11)
Considering our problem, we, as in [8], define
@@
(a vector version of (3.11)) which clearly implies
@@
and then prove in Appendix B that functions (3.5)–(3.8) satisfy (2.7) and (2.8). Moreover, these functions are solutions of all differential equations of the ALH (equations (2.13)–(2.16) included), provided the variables of the hierarchy, $z_{1},z_{2},$..., are introduced in accordance with the definition (3.12). Returning to our problem, this means that the $(x,t)$-dependency is governed by
@@
where $\langle a_{0}|=\left(a_{01},...,a_{0N}\right)$ and $\langle b_{0}|=\left(b_{01},...,b_{0N}\right)$ are arbitrary constant NN-rows,
@@
and $\mathsf{D}_{\alpha}$, $\mathsf{D}_{\beta}$, $\mathsf{D}_{a}$, $\mathsf{D}_{b}$ are diagonal matrices with elements $\alpha_{j}$, $\beta_{j}$, $a_{0j}$, $b_{0j}$ (j=1,,Nj=1,...,N) correspondingly, one can present $\mathsf{A}$ and $\mathsf{B}$ as
$\mathsf{A}=\mathsf{D}_{\alpha}\mathsf{Y}\mathsf{D}_{\beta}^{-1},\quad\mathsf{B}=\mathsf{D}_{\beta}\mathsf{Y}^{*}\mathsf{D}_{\alpha}^{-1}$ (3.25)
@@
$\mathsf{Y}=\mathsf{C}\mathsf{D}_{\beta}\mathsf{D}_{a}\mathsf{E}_{A}.$ (3.26)
which leads to
$\mathsf{F}=\mathsf{D}_{\beta}\left(\mathsf{1}+\mathsf{Y}^{*}\mathsf{Y}\right)^{-1}\mathsf{D}_{\beta}^{-1},\qquad\mathsf{G}=\mathsf{D}_{\alpha}\left(\mathsf{1}+\mathsf{Y}\mathsf{Y}^{*}\right)^{-1}\mathsf{D}_{\alpha}^{-1}.$ (3.27)
@@
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\langle\gamma|\mathsf{\Omega}|1\rangle,\quad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}=1+\langle\gamma|\mathsf{\Omega}\mathsf{Y}^{*}|1\rangle$ (3.28)
@@
$\mathsf{\Omega}=\mathsf{Y}\left(\mathsf{1}+\mathsf{Y}^{*}\mathsf{Y}\right)^{-1},$ (3.29)
the constant row $\langle\gamma|$ is given by
$\langle\gamma|=\langle 1|\left(\mathsf{C}\mathsf{L}^{*}\right)^{-1}$ (3.30)
@@
In a similar way, the ansatz for $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}$, given by (3.6) and (3.8), rewritten in terms of $\mathsf{Y}$, $\mathsf{\Omega}$ and $\langle\gamma|$ leads to
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}=\langle\gamma^{*}|\mathsf{\Omega}^{*}|1\rangle,\quad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=1+\langle\gamma^{*}|\mathsf{\Omega}^{*}\mathsf{Y}|1\rangle$ (3.32)
which demonstrates that $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}=\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}^{\!*}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}^{\!*}$.
The dependence on $t$ and $x$ is described by the matrix $\mathsf{E}$,
$\mathsf{E}=\mathsf{D}_{\beta}\mathsf{D}_{a}\mathsf{E}_{A}$ (3.33)
which can be written as
$\mathsf{E}=\mathop{\mbox{diag}}\left(\;e^{f_{k}+i\varphi_{k}}\;\right)_{k=1,...,N}$ (3.34)
with
$\begin{array}[]{lcl}f_{k}(t,x)&=&\nu_{k}x+2\mu_{k}\nu_{k}t+f_{0k},\\[5.69054pt] \varphi_{k}(t,x)&=&-\mu_{k}x+(\nu_{k}^{2}-\mu_{k}^{2})t+\varphi_{0k}\end{array}$ (3.35)
where $\mu_{k}$ and $\nu_{k}$ are defined by
$\mu_{k}+i\nu_{k}=L_{k}\left/|L_{k}|^{2}\right$. (3.36)
and $f_{0k}$, $\varphi_{0k}$ are arbitrary constants.
@@
$\boldsymbol{\psi}=\mathsf{U}\left(\begin{array}[]{c}\langle\gamma|\mathsf{\Omega}|1\rangle,\quad\\ 1+\langle\gamma|\mathsf{\Omega}\mathsf{Y}^{*}|1\rangle\end{array}\right)$ (3.37)
where $\mathsf{U}$ is an arbitrary constant unitary matrix,
$\mathsf{\Omega}=\mathsf{Y}\left(\mathsf{1}+\mathsf{Y}^{*}\mathsf{Y}\right)^{-1},\quad\mathsf{Y}=\mathsf{C}\mathsf{E},$ (3.38)
@@
with $\langle\gamma|$, $\mathsf{C}$ and $\mathsf{E}$ defined in (3.30), (3.24) and (3.34), solve the Euler–Lagrange equations (1.3). Elements of the matrix $\mathsf{U}$ together with the $4N$ constants $\mathrm{Re}\,L_{k}$, $\mathrm{Im}\,L_{k}$, $f_{0k}$ and $\varphi_{0k}$ are arbitrary parameters that determine the properties of the $N$-soliton solution.
To get some insight into the structure of obtained solutions, let us consider the one-soliton solution.
In this case the matrices $\mathsf{L}$, $\mathsf{C}$ and the rows $\langle\gamma|$ become just scalars,
$\mathsf{L}\to L_{1},\quad\mathsf{C}\to\frac{1}{2i|L_{1}|\sin\theta},\quad\langle\gamma|\to e^{2i\theta}-1$ (3.39)
where $\theta=\arg L_{1}$, or
$\theta=\tan^{-1}(\nu_{1}/\mu_{1}).$ (3.40)
Modifying slightly the functions $f_{1}$ and $\varphi_{1}$ that appear in (3.34), i.e. introducing the new ones, $f$ and $\varphi$, defined by
$f(t,x)=f_{1}(t,x)+\ln\frac{\mu_{1}^{2}+\nu_{1}^{2}}{2|\nu_{1}|},$ (3.41)
$\varphi(t,x)=\varphi_{1}(t,x)-\frac{\pi}{2}\mathop{\mbox{sign}}\nu_{1}$ (3.42)
one can present $\mathsf{Y}$ as $\mathsf{\Omega}$ as
$\mathsf{Y}=\exp(f+i\varphi),\quad\mathsf{\Omega}=\frac{1}{2}\exp(i\varphi)\mathop{\mbox{sech}}f.$ (3.43)
Choosing $\mathsf{U}=\mbox{diag}\left(e^{-i\theta},e^{-i\theta}\right)$ one can obtain the following expressions for the components of the vector $\boldsymbol{\psi}=\left(\psi_{1},\psi_{2}\right)^{\scriptscriptstyle T}$:
$\psi_{1}=i\sin\theta\,\exp(i\varphi)\mathop{\mbox{sech}}{f},$ (3.44)
$\psi_{2}=\cos\theta+i\sin\theta\,\tanh f.$ (3.45)
with
$f(t,x)=\nu_{1}x+2\mu_{1}\nu_{1}t+f_{0},$ (3.46)
$\varphi(t,x)=-\mu_{1}x+(\nu_{1}^{2}-\mu_{1}^{2})t+\varphi_{0}.$ (3.47)
It is easy to see that $\psi_{1}$ and $\psi_{2}$ are the NLSE-solitons of different type: $\psi_{1}$ is a so-called bright soliton, vanishing as $x\to\pm\infty$ (with $t$ being fixed), while $\psi_{2}$ is a dark soliton ($\lim_{|x|\to\infty}\psi_{2}=e^{\pm i\theta}$). However, this does not mean that the same is true for any solution (even for the one-soliton one), because for an arbitrary $\mathsf{U}$, all components of the general $N$-soliton solution are mixtures of dark and bright solitons.
Another useful information that can be obtained form (3.44)–(3.47) is the fact that all physically important characteristics of the soliton, its amplitude ($\sin\theta$), velocity ($2\mu_{1}$) and the scale ($1/\nu_{1}$), are determined by the choice of $L_{1}$. The same is true in the multi-soliton case. Although a $N$-soliton solution of a nonlinear equation is surely not a sum of $N$ solitons (only in the asymptotic regions, $t\to\pm\infty$ it can be viewed as such), it is possible to describe its structure qualitatively in terms of single solitons whose amplitude, velocity and scale depend on on the elements of the matrix $\mathsf{L}$, $L_{k}$, while other parameters, $\alpha_{k}$, $\beta_{k}$, $a_{0k}$ and $b_{0k}$, which were ‘absorbed’ into the matrix $\mathsf{E}$ (3.34), i.e. replaced by $f_{0k}$ and $\varphi_{0k}$, determine their relative position and phasing. The role of the matrix $\mathsf{U}$ is mostly ‘mixing’ of different solitons.
4 Vector NLSE with gradient nonlinearity.
In this section we consider an example of application of the results, obtained in this paper for a rather abstract system, to a more physical problem.
Let us return to the system (2.21), (2.22) and study the behavior of the function $s$ defined as
$s=\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}_{x}=-\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}$ (4.1)
(the last equality stems from the fact $\partial_{x}(\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q})=0$). Calculating the derivative of $s$ using both of the equations in (4.1) and expressing $\boldsymbol{q}_{xx}$ and $\boldsymbol{r}^{\scriptscriptstyle T}_{xx}$ from (2.21) and (2.22), one can easily obtain
$s_{x}$ $=$ $-i\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}_{t}+\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x}-\lambda$ (4.2)
$=$ $-i\boldsymbol{r}^{\scriptscriptstyle T}_{t}\boldsymbol{q}-\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x}+\lambda$ (4.3)
which, after summation of both expressions, leads to
$s_{x}=-(i/2)\,\partial_{t}(\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q})=0$ (4.4)
which means that $s=s(t)$.
In a similar way, one can calculate the time derivative of $s$ which leads to
$is_{t}=\partial_{x}\left(2\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x}-\lambda\right).$ (4.5)
This implies that
$\lambda=\lambda_{0}+2\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x}$ (4.6)
where
$\lambda_{0}=-is^{\prime}(t)x+s_{1}(t)$ (4.7)
and $s_{1}$ is another functions not depending on $x$. Thus, equations (2.21) and (2.22) can be rewritten as
$i\boldsymbol{q}_{t}+\boldsymbol{q}_{xx}+(\lambda_{0}+2\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x})\boldsymbol{q}=0,$ (4.8)
$i\boldsymbol{r}^{\scriptscriptstyle T}_{t}+\boldsymbol{r}^{\scriptscriptstyle T}_{xx}+(\lambda_{0}+2\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x})\boldsymbol{r}^{\scriptscriptstyle T}=0$ (4.9)
In the case of the soliton solutions presented in the previous section, it can be shown, by calculating $\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x}$ using equations (2.18) and comparing the result with the expression for $\lambda$ in (2.23), that $\lambda_{0}=0$. Thus, as a byproduct of the calculations presented in this paper, we obtain the following result.
Proposition 4.1
NN-soliton solutions described in proposition 3.2 are, at the same time, solutions of the NLSE with gradient nonlinearity
$i\boldsymbol{\psi}_{t}+\boldsymbol{\psi}_{xx}+2\,\boldsymbol{\psi}^{\dagger}_{x}\boldsymbol{\psi}_{x}\,\boldsymbol{\psi}=0.$ (4.10)
@@
A more interesting question is whether we can tackle with the approach of this paper the case on quadratic restrictions other than (1.2), for example, ones given by